Two medium eggs and three cups of ice cream contain 701 milligrams of cholesterol. One medium egg and one cup of ice cream exceed the suggested daily cholesterol intake of 300 milligrams by 25 milligrams. Determine the cholesterol content in each item.
Answer by stanbon(75887)

Two medium eggs and three cups of ice cream contain 701 milligrams of cholesterol. One medium egg and one cup of ice cream exceed the suggested daily cholesterol intake of 300 milligrams by 25 milligrams. Determine the cholesterol content in each item.
------------------
EQUATIONS:
1st: 2m + 3c = 701
2nd: 1m + 1c = 325
------------------
Multiply 2nd by 2 to get:
3rd: 2m + 2c = 650
--------------
Substract 1st from 3rd to get:
-c = - 51
c = 51 (1 cup of ice cream has 51 mg of cholesterol)
---------------
Substitute into m+c 325 to solve for m:
m + 51 = 325
m = 274 mg (one medium egg has 274 mg of cholesterol)
==========
Cheers,
Stan H.